2KClO3 → 2KCl + 3O2
What is the % yield of the above reaction if only 45 g of oxygen is produced from 122 g of potassium chlorate?
1
Expert's answer
2020-06-04T10:37:41-0400

First the theoretical yield of O2 should be determined.

n (KClO3) = 122g / 122g/mol = 1 mol

According to the reaction, n (O2) = 3/2 x n (KClO3) = 1.5 mol

m (O2) = 1.5 mol x 32 g/mol = 48 g - theoretical yield.

%yield = (actual yield / theoretical yield) x 100% = 45/48 x 100% = 93.75%.

Answer: 93.75%